OpenStudy (anonymous):

The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show your work

OpenStudy (anonymous):

The formula would be \[SA = 2 \pi r ^{2} + 2 \pi r h\] so you would do \[2\pi(3.5)^{2}+2\pi(3.5)(14)\] and whatever that is, is your answer!

OpenStudy (anonymous):

ok thnkz :D

OpenStudy (anonymous):

384.65^3